# A LLM-Powered Automatic Grading Framework with Human-Level Guidelines Optimization

### v1

@@ Abstract.
Open-ended short-answer questions (SAGs) have been widely recognized as a powerful tool for providing deeper insights into learners’ responses in the context of learning analytics (LA). However, SAGs often present challenges in practice due to the high grading workload and concerns about inconsistent assessments. With recent advancements in natural language processing (NLP), automatic short-answer grading (ASAG) offers a promising solution to these challenges. Despite this, current ASAG algorithms are often limited in generalizability and tend to be tailored to specific questions. In this paper, we propose a unified multi-agent ASAG framework, GradeOpt, which leverages large language models (LLMs) as graders for SAGs. More importantly, GradeOpt incorporates two additional LLM-based agents—the reflector and the refiner—into the multi-agent system. This enables GradeOpt to automatically optimize the original grading guidelines by performing self-reflection on its errors. Through experiments on a challenging ASAG task, namely the grading of pedagogical content knowledge (PCK) and content knowledge (CK) questions, GradeOpt demonstrates superior performance in grading accuracy and behavior alignment with human graders compared to representative baselines. Finally, comprehensive ablation studies confirm the effectiveness of the individual components designed in GradeOpt.
Large Language Models, Automatic Grading, Mathematics Education, Professional Development
1. Introduction
Learning analytics (LA), supported by data processing and performance prediction, presents enormous educational opportunities to enhance learning quality through the provision of effective feedback and actions (Bienkowski et al., 2012). Meanwhile, providing timely and precise assessments to student assignments and examinations is vital to LA due to the significance of performance measurement in the learning process (Süzen et al., 2020). Traditionally, close-ended multiple-choice questions (MCQs), which asks students to select correct answer from distracting options, dominated LA studies. This is due to their convenience in the grading process (Mohler et al., 2011; Burrows et al., 2015). However, open-ended short-answer questions (SAQs), which asks for not only final answer but explanations, often provide deeper insights into students’ answering rationale and knowledge concepts. This is because they are known to elicit the thinking path that describes how a student arrives at their conclusion (Leacock and Chodorow, 2003). Unfortunately, grading open-ended textual answers is tedious as the workloads are significantly increased, and instructors need to examine each response and provide feedback (Roy et al., 2015). More importantly, the inconsistent or unfair assessments, caused by diverged interpretations, biases, or mistakes create another challenge to SAQs grading in practice (Süzen et al., 2020). To mitigate these issues and provide timely and consistent evaluation, automatic short-answer grading (ASAG) (Burrows et al., 2015) systems have become appealing.
@@ 1. Introduction
ASAG, which can be traced back to the 1960s, has bloomed in recent years due to advancements in natural language processing (NLP) (Leacock and Chodorow, 2003; Xie et al., 2024). Early ASAG systems often used pattern-matching techniques and hand-crafted features (Leacock and Chodorow, 2003). Thus, those systems required intensive human labor to build and were limited to a few specific grading tasks. The rise of deep learning (DL) has lessened the amount of burdensome feature designs needed for early ASAG systems. DL provides an end-to-end solution that automatically learns to output grading scores from a large number of graded answer samples (Hassan et al., 2018). Due to the strong data-fitting capability of DL models, DL-based ASAG systems are able to be extended to different tasks if a large number of annotated samples are available. However, when the annotated sample size is limited, DL-based ASAG systems often face serious over-fitting issues. Beyond that, as DL is a black-box model whose results lack interpretation, the application of DL-based ASAG systems is still limited (Condor and Pardos, 2024).
@@
The emergence of pre-trained language models (PLMs) and the more advanced Large Language Models (LLMs) have recently revolutionized the design of ASAG systems due to their human-like language ability and human-interpretable intermediate textual results. Therefore, many recent studies have attempted to build ASAG systems with LLMs. Promising results have been demonstrated that using fine-tuning (Latif and Zhai, 2023) and prompting techniques such as Chain-of-Thought (CoT) (Cohn et al., 2024) and in-context learning (Lee et al., 2024). Yet these recent techniques are still limited due to LLMs’ inherent limitations such as sensitivity to prompts, context window restriction, etc., making the complex ASAG task challenging for the LLM grader. In reality, accurate, standardized, and unambiguous guidelines are critical to help human graders formulate a precise interpretation of scoring criteria. For LLM-based ASAG systems, those guidelines also serve as the principal instructions. They teach LLMs to perform the grading task following a similar standard as human graders. However, using guidelines composed by pedagogical experts directly for LLMs is sub-optimal since the general-purposed LLMs lack domain-specific knowledge and can misinterpret the guidelines (Leacock and Chodorow, 2003). Meanwhile, LLMs are often sensitive to various facets of prompts (Jiang et al., 2020) where minor changes could lead to great differences in LLM’s performance. Optimizing the guidelines manually for LLMs can further take a lot of trial and error. Thus, recent works propose to conduct guideline modification with LLMs to offload human burden (Cohn et al., 2024). While the modified guidelines yield performance improvement, the prompt search space in these methods is relatively limited. Because of this, the modified guidelines are not necessarily optimal. Additionally, abundant human efforts such as timely feedback or a large amount of labeling are required. Therefore, methods to optimize grading guidelines automatically and effectively are still desired.
@@
In this paper, we propose a unified multi-agent ASAG framework that can automatically optimize grading guidelines. Specifically, it utilizes an iterative reflection mechanism to generate a task prompt (guideline) that can capture learners’ thinking and knowledge from a small dataset of short answers. To do this, we innovatively propose to leverage prompt optimization in ASAG by viewing grading guideline refinement as an optimization problem that aims for maximal accuracy. To implement the backbone algorithm, we take inspiration from APO (Pryzant et al., 2023) and propose novel techniques such as misconfidence-based selection, iterative optimization, and log-probability-based robustness to enhance the ASAG system’s stability in making accurate and trustable score predictions to unseen datasets. To minimize human labelling effort, our mechanism intelligently selects short answer samples that can lead to the optimal performance of guideline enhancement. Additionally, our framework serves as a solution to test questions of a variety of complexity, providing interpretable assessments to each learning goal that contributes to the overall scoring. To validate the effectiveness of the framework, we conducted experiments with data collected from teachers who answered a set of questions designed to capture their content knowledge (CK) and pedagogical content knowledge (PCK) of mathematics (Copur-Gencturk et al., 2023). The experimental results demonstrate that both accuracy and alignment increase as compared to the representative baselines. In addition, our further analysis showcases steady improvement of test accuracy throughout iterations, demonstrating effective and continuous enhancement of the learned guidelines. To the best of our knowledge, we are the first to apply prompt optimization in ASAG by optimizing grading guidelines similar to how an optimal task prompt is generated. We believe that refining the grading guidelines through a multi-agent reflective mechanism can help the great potential of LLMs in LA by better assisting learners with detailed, accurate assessment while significantly reducing educators’ scoring burden.
2. Related Work
2.1. Automatic Short Answer Grading
Automatic Short Answer Grading (ASAG) is often treated as a text classification or regression problem in NLP studies. Here we mainly focus on classification due to its relevancy to our setting. Traditional ASAG models mainly rely on text similarity and employ classic ML classifiers. They use lexical features such as bag-of-words (BOW) (Mohler et al., 2011) and TF-IDF (Gobbo et al., 2023), or syntactic features indicating the structure of sentences (Leacock and Chodorow, 2003). However, these methods require significant manual design, which makes them hard to be applied to new datasets. To reduce the burden of feature engineering, Deep Neural Networks (DNNs) such as Long-Short-Term-Memory (LSTM) are utilized (Hassan et al., 2018), which produce superior results but suffer from limited generalizability. Pre-trained BERT-based models provide enhanced versatility through transfer learning including on ASAG datasets (Camus and Filighera, 2020). To further enhance grading accuracy, researchers have made attempts to ensemble BERT with statistics-based methods  (Erickson et al., 2020) and data augmentation (Lun et al., 2020). LLMs are increasingly utilized in ASAG and similar assessment tasks (Cohn et al., 2024; Xie et al., 2024). However, their prompts are mostly always manually-crafted and thus are unable to properly adapt to new datasets. To solve this issue, several works have shifted attention to assisting educators with guideline creation (Xie et al., 2024; Cohn et al., 2024).
2.2. LLM Prompt Optimization and Reflection
Prompts are critical to the success of LLMs (Zhou et al., 2023). To tailor LLMs to challenging tasks, manually crafted prompts are adopted to enhance the performance (Wei et al., 2023). To automate the generation and optimization of prompts, prompt optimization emerges as a promising method for input prompt refinement. Using these techniques, LLMs have demonstrated superior performance in many down-stream tasks, particularly in instruction following and reasoning (Pryzant et al., 2023; Zhou et al., 2023; Yang et al., 2024). However, such automatic methods are risky when directly applied to ASAG tasks considering the limitations of LLMs such as hallucination (Huang et al., 2023) and misalignment (Kenton et al., 2021). To enhance both accuracy and trustworthiness, we adopt the idea of state-of-the-art prompt optimization APO (Pryzant et al., 2023) and implement novel techniques for reliability.
Similar to how humans gather knowledge from failures, experience-and-reflect (Pan et al., 2023) is an important technique for improving LLMs’ alignment with task specifications. By reflection, LLMs learn through failure, which enriches its knowledge base and provides valuable reference in similar scenarios. Self-reflection has demonstrated promising results in improving LLM reasoning (Shinn et al., 2023; Madaan et al., 2023). However, LLMs’ reflection ability is relatively limited when it comes to self-correction without human feedback or true labels (Huang et al., 2024). A recent work (Tyen et al., 2024) divides the task of self-correction into two steps: mistake finding and output correction. They empirically show that while LLMs struggle to find errors, their correction ability is robust when given ground-truth labels. This provides grounding support for our proposed framework due to the similar use of true labels in guiding LLM reflection.
3. Problem Statement
@@
4. Method
@@ 4.1. An Overview
Figure 1. An Illustration of the Proposed Framework.
4.2. Training Stage
@@
Before diving into the details of this stage, we first give a brief introduction to the three key components of a common grading guideline: Question Stem ($G_{qs}$), Key Concept ($G_{kc}$) and Scoring Rubric ($G_{sr}$). Specifically, $G_{qs}$ contains the complete question contents, $G_{kc}$ describes the test knowledge concepts, and $G_{sr}$ is the operational guidance instructing human graders how to score responses. As we previously mentioned, directly using $G_{0}=\{G_{qs}||G_{kc}||G_{sr}\}$ as grading guideline for Grader is sub-optimal since the human-based scoring rubrics commonly lack detailed explanations to some concepts. As a result, LLM-based grading methods could provide ambiguous judgments. To solve this issue, GradeOpt focuses on optimizing $G$ by appending new Adaption Rules ($G_{ar}$) that provides the detailed explanations regarding reflections from failed predictions and identified errors. In Figure  2, we present an example of optimized grading guideline $G_{opt}$. Specifically, when given the expert-designed input containing ”Task Description”, ”Question Stem”, ”Key Concept” and ”Scoring Rubrics”, GradeOpt automatically generates the additional descriptions in ”Adaptation Rules”. These new rules help describe how to assign a grade based on answer patterns and details.
@@
Task Description In this task, you perform the task of assessing teachers’ knowledge of students’ mathematical thinking by grading their responses to a pedagogical content knowledge question. Question Stem Based on the student’s work, what is student likely to understand about the relationship between the length of the shadow and the height of the object? Key Concept Teachers should infer that the student possibly understands that there is a proportional relationship between the height of an object and the length of its shadow. However, because the concept of halving/doubling is natural to students, it is unclear if the student understands the relationship between object height and shadow length is proportional or if they understand equivalent ratios. Scoring Rubrics - Award 0 points if the response does not address the key concept … - Award 1 point if the response includes an accurate mention or implicit understanding of the key concept, even if it is contextually incomplete or unclear … - Award 2 points if the response offers a clear and explicit analysis of the proportional relationship between the objects, demonstrating that the ratio/proportion is consistent … Adaptation Rules 1. Mention of Key Concept (Proportionality/Equivalent Ratios): - Rule: Award 1 point if the teacher response includes any accurate mention or implicit understanding of the Key Concept (proportionality/equivalent ratios), even if it lacks detailed analysis or evidence from the student’s work. - Example 1: If the response mentions that the student understands the relevance of using equivalent ratios, this should earn 1 point even if the analysis is not detailed. - Example 3: If the response mentions the concept of equivalent ratios and provides an example (1:2 ratio), this should earn 1 point despite the explanation being somewhat unclear and lacking depth. - Example 4: If the response mentioned proportionality and suggested comparing the sides, this should earn 1 point even if it lacks detailed analysis or evidence from the student’s work.
2. Clear and Explicit Analysis of Proportional Relationship: - Rule: Award 2 points if the teacher response offers a clear and explicit analysis of the proportional relationship between the objects, demonstrating that the ratio/proportion is consistent, even if it does not provide specific evidence from the student’s work. - Example 2: If the response provides a clear and explicit analysis of the proportional relationship between the lengths of shadows and the heights of objects, explaining the reasoning behind this relationship (constant angle of sunlight), this should earn 2 points.
3. Detailed Analysis and Specific Evidence: - Rule: Award 2 points if the response includes both an explicit analysis of the student’s understanding of proportionality/equivalent ratios (Key Concept) and clear, specific evidence from the student’s work supporting this analysis. - Example: If a response explicitly analyzes the student’s understanding of proportionality and cites specific parts of the student’s work (e.g., ”Student A’s work shows that they understand the flagpole’s shadow is half its height, and they apply this to the lighthouse”), it should earn 2 points.
Figure 2. An example of the optimized guidelines, $G^{*}$.
4.2.1. Agent Configurations

### v2

@@ Abstract
Open-text responses provide researchers and educators with rich, nuanced insights that multiple-choice questions cannot capture. When reliably assessed, such responses have the potential to enhance teaching and learning. However, scaling and consistently capturing these nuances remain significant challenges, limiting the widespread use of open-text questions in educational research and assessments. In this paper, we introduce and evaluate GradeOpt, a unified multi-agent automatic short-answer grading (ASAG) framework that leverages large language models (LLMs) as graders for short-answer responses. More importantly, GradeOpt incorporates two additional LLM-based agents—the reflector and the refiner—into the multi-agent system. This enables GradeOpt to automatically optimize the original grading guidelines by performing self-reflection on its errors. To assess GradeOpt’s effectiveness, we conducted experiments on two representative ASAG datasets, which include items designed to capture key aspects of teachers’ pedagogical knowledge and students’ learning progress. Our results demonstrate that GradeOpt consistently outperforms representative baselines in both grading accuracy and alignment with human evaluators across different knowledge domains. Finally, comprehensive ablation studies validate the contributions of GradeOpt’s individual components, confirming their impact on overall performance.
keywords:
Large Language Models, Automatic Grading, Mathematics Education, Teacher Knowledge, Assessments, Large-Scale Testing, Mathematical Knowledge for Teaching, Physical Science, Learning Assessments, Content Knowledge, Pedagogical Content Knowledge
titlenote: Both authors contribute equally.
1 Introduction
@@
Accurate evaluation of assignments and examinations in a timely manner is vital to learning due to the significance of performance measurement in the learning process suzen2020textmining . Traditionally, multiple-choice questions (MCQs), which asks students to select the correct answer from distracting options, dominated learning assessment studies. While this approach makes data available promptly  (mohler2011learning, ; burrows2015eras, ), it falls short in proving insights into learners’ thinking. Open-ended short-answer questions (SAQs) can provide deeper insights into students’ answering rationale and knowledge concepts. This is because they are known to elicit the thinking path that describes how a student arrives at their conclusion (leacock2003c-rater, ). Unfortunately, grading open-ended textual answers is tedious as substantial resources and time are needed to train raters to accurately and consistently code responses (roy2015perspective, ). More importantly, the inconsistent or unfair assessments, caused by diverged interpretations, biases, or mistakes create another challenge to SAQs grading in practice (suzen2020textmining, ). To mitigate these issues and provide timely and consistent evaluation, automatic short-answer grading (ASAG) burrows2015eras systems have become appealing. ASAG, which can be traced back to the 1960s, has bloomed in recent years due to advancements in natural language processing (NLP) leacock2003c-rater ; xie2024grade . Early ASAG systems often used pattern-matching techniques and hand-crafted features leacock2003c-rater . Thus, those systems required intensive human labor to build and were limited to a few specific grading tasks. The rise of deep learning (DL) has lessened the amount of burdensome feature designs needed for early ASAG systems. DL provides an end-to-end solution that automatically learns to output grading scores from a large number of graded answer samples hassan2018automatic . Due to the strong data-fitting capability of DL models, DL-based ASAG systems are able to be extended to different tasks if a large number of annotated samples are available. However, when the annotated sample size is limited, DL-based ASAG systems often face serious over-fitting issues. Beyond that, as DL is a black-box model whose results lack interpretation, the application of DL-based ASAG systems is still limited condor2024explainable .
@@
The emergence of pre-trained language models (PLMs) and the more advanced Large Language Models (LLMs) have recently revolutionized the design of ASAG systems due to their human-like language ability and human-interpretable intermediate textual results. Therefore, many recent studies have attempted to build ASAG systems with LLMs. Promising results have been demonstrated that using fine-tuning latif2023finetuning and prompting techniques such as Chain-of-Thought (CoT) cohn2024science and in-context learning lee2024applyingllm . Yet these recent techniques are still limited due to LLMs’ inherent limitations such as sensitivity to prompts, context window restriction, etc., making the complex ASAG task challenging for the LLM grader. In reality, accurate, standardized, and unambiguous guidelines are critical to help human graders formulate a precise interpretation of scoring criteria. For LLM-based ASAG systems, those guidelines also serve as the principal instructions. They teach LLMs to perform the grading task following a similar standard as human graders. However, using guidelines composed by pedagogical experts directly for LLMs is sub-optimal since the general-purposed LLMs lack domain-specific knowledge and can misinterpret the guidelines (leacock2003c-rater, ). Meanwhile, LLMs are often sensitive to various facets of prompts (jiang2020knowlmsknow, ) where minor changes could lead to great differences in LLM’s performance. Optimizing the guidelines manually for LLMs can further take a lot of trial and error. Thus, recent works propose to conduct guideline modification with LLMs to offload human burden cohn2024science . While the modified guidelines yield performance improvement, the prompt search space in these methods is relatively limited. Because of this, the modified guidelines are not necessarily optimal. Additionally, abundant human efforts such as timely feedback or a large amount of labeling are required. Therefore, methods to optimize grading guidelines automatically and effectively are still desired.
@@
In this paper, we propose a unified multi-agent ASAG framework that automatically optimizes grading guidelines. Specifically, it employs an iterative reflection mechanism to generate task prompts (guidelines) that effectively capture learners’ thinking and knowledge from a small dataset of short answers. To achieve this, we innovatively introduce prompt optimization in ASAG, framing grading guideline refinement as an optimization problem aimed at maximizing accuracy. Inspired by APO pryzant2023gradient , we develop novel techniques such as misconfidence-based selection, iterative optimization, and log-probability-based robustness to enhance the framework’s stability in producing accurate and trustworthy score predictions on unseen datasets. To minimize human labeling effort, our mechanism intelligently selects short-answer samples that contribute to optimal guideline refinement. Additionally, the framework supports assessments across varying levels of complexity, offering interpretable evaluations for each learning objective while improving overall scoring accuracy. To validate our approach, we conducted experiments on two real-world grading datasets. The first dataset comprises responses from high school students within a physical sciences curriculum, while the second consists of a national sample of teachers answering questions designed to assess content-specific knowledge required for teaching (copur2022mathematics, ). Experimental results demonstrate that GradeOpt outperforms representative baselines in both accuracy and alignment. Further analysis highlights consistent improvements in test accuracy across iterations, showcasing the framework’s ability to continuously enhance grading guidelines. To the best of our knowledge, we are the first to apply prompt optimization in ASAG by refining grading guidelines akin to generating an optimal task prompt. We believe that our multi-agent reflective mechanism can unlock the full potential of LLMs in learning analytics by providing detailed and accurate assessments while significantly reducing educators’ grading workload.
@@
Illustration of the proposed framework Refer to caption
Figure 1: An Illustration of the proposed framework.
2 Related Work
2.1 Automatic Short Answer Grading
Automatic Short Answer Grading (ASAG) is often treated as a text classification or regression problem in NLP studies. Here we mainly focus on classification due to its relevancy to our setting. Traditional ASAG models mainly rely on text similarity and employ classic ML classifiers. They use lexical features such as bag-of-words (BOW) mohler2011learning and TF-IDF del2023gradeaid , or syntactic features indicating the structure of sentences leacock2003c-rater . However, these methods require significant manual design, which makes them hard to be applied to new datasets. To reduce the burden of feature engineering, Deep Neural Networks (DNNs) such as Long-Short-Term-Memory (LSTM) are utilized hassan2018automatic , which produce superior results but suffer from limited generalizability. Pre-trained BERT-based models provide enhanced versatility through transfer learning including on ASAG datasets camus2020investigating . To further enhance grading accuracy, researchers have made attempts to ensemble BERT with statistics-based methods  erickson2020automated and data augmentation lun2020multiple . LLMs are increasingly utilized in ASAG and similar assessment tasks yang2024content ; cohn2024science ; chu2025enhancing . However, their prompts are mostly always manually-crafted and thus are unable to properly adapt to new datasets. To solve this issue, several works have shifted attention to assisting educators with guideline creation xie2024grade ; cohn2024science .
2.2 LLM Prompt Optimization and Reflection
Prompts are critical to the success of LLMs zhou2023ape . To tailor LLMs to challenging tasks, manually crafted prompts are adopted to enhance the performance wei2023cot . To automate the generation and optimization of prompts, prompt optimization emerges as a promising method for input prompt refinement.
Using these techniques, LLMs have demonstrated superior performance in many down-stream tasks, particularly in instruction following and reasoning pryzant2023gradient ; zhou2023ape ; yang2024large . However, such automatic methods are risky when directly applied to ASAG tasks considering the limitations of LLMs such as hallucination huang2023surveyhallucinationllm and misalignment kenton2021alignment . To enhance both accuracy and trustworthiness, we adopt the idea of state-of-the-art prompt optimization APO pryzant2023gradient and implement novel techniques for reliability. Similar to how humans gather knowledge from failures, experience-and-reflect pan2023automatically is an important technique for improving LLMs’ alignment with task specifications. By reflection, LLMs learn through failure, which enriches its knowledge base and provides valuable reference in similar scenarios. Self-reflection has demonstrated promising results in improving LLM reasoning shinn2023reflexion ; madaan2023selfrefine . However, LLMs’ reflection ability is relatively limited when it comes to self-correction without human feedback or true labels huang2024largelanguagemodelsselfcorrect . A recent work tyen2024llmscannot divides the task of self-correction into two steps: mistake finding and output correction. They empirically show that while LLMs struggle to find errors, their correction ability is robust when given ground-truth labels. This provides grounding support for our proposed framework due to the similar use of true labels in guiding LLM reflection.
3 Problem Statement
[…]
4 Method
@@ 4.1 An Overview
Task Description: You are GradeGPT. You assess teachers’ knowledge of students’ mathematical thinking by grading their responses to a pedagogical content knowledge question.
Question Stem: Based on the student’s work, what is student likely to understand about the relationship between the length of the shadow and the height of the object?
Key Concept: Teachers should infer that the student possibly understands that there is a proportional relationship between the height of an object and the length of its shadow. However, because the concept of halving/doubling is natural to students, it is unclear if the student understands the relationship between object height and shadow length is proportional or if they understand equivalent ratios.
Scoring Rubrics: - Award 0 points if it does not address key concept … - Award 1 point if the response includes an accurate mention or implicit understanding of the key concept … - Award 2 points if the response offers a clear and explicit analysis of the proportional relationship between the objects …
Adaptation Rules: 1. Mention of Key Concept: - Rule: Award 1 point if the teacher response includes any accurate mention or implicit understanding of the Key Concept (proportionality/equivalent ratios), even if it lacks detailed analysis or evidence from the student’s work. - Example 1: If the response mentions that the student understands the relevance of using equivalent ratios, this should earn 1 point even if the analysis is not detailed.
Figure 2: An example of the optimized guidelines, $G^{*}$.
4.2 Training Stage
@@
Before diving into the details of this stage, we first give a brief introduction to the three key components of a common grading guideline: Question Stem ($G_{qs}$), Key Concept ($G_{kc}$) and Scoring Rubric ($G_{sr}$). Specifically, $G_{qs}$ contains the complete question contents, $G_{kc}$ describes the test knowledge concepts, and $G_{sr}$ is the operational guidance instructing human graders how to score responses. As we previously mentioned, directly using $G_{0}=\{G_{qs}||G_{kc}||G_{sr}\}$ as grading guideline for Grader is sub-optimal since the human-based scoring rubrics commonly lack detailed explanations to some concepts. As a result, LLM-based grading methods could provide ambiguous judgments. To solve this issue, GradeOpt focuses on optimizing $G$ by appending new Adaption Rules ($G_{ar}$) that provides the detailed explanations regarding reflections from failed predictions and identified errors. In Figure 2, we present an example of optimized grading guideline $G_{opt}$. Specifically, when given the expert-designed input containing “Task Description", “Question Stem", “Key Concept" and “Scoring Rubrics", GradeOpt automatically generates the additional descriptions in “Adaptation Rules". These new rules help describe how to assign a grade based on answer patterns and details.
@@
4.2.1 Agent Configurations